mathslover (mathslover):

If n is positive integer, prove that the integral part of \((5\sqrt{5} + 11)^{2n+1} \) is an even number.

OpenStudy (dan815):

what does it mean integral part

mathslover (mathslover):

1 + sqrt(3) <--- integral part is 1 2/3 + 1 <--- integral part is 1

OpenStudy (dan815):

okay i see

OpenStudy (dan815):

well u can look at the binomial expansion

OpenStudy (dan815):

odd^odd=odd odd^even=odd

OpenStudy (dan815):

and the addition of 2 odds = even

OpenStudy (dan815):

therefore the integral part can only be some summation of odd^odd and odd^even there will be a total of 2n+2 of these addition going from 0 to 2n+1

OpenStudy (dan815):

therefore an even number of addition of these odd^2 and odd^even so an even number of odd adition gives you an even number in the end

OpenStudy (dan815):

and a simple side note*, is that there is always a even number coming out of the choosing formula except for 0! and 1! and 2n+1 choose 2n+1, which is just 1, or u can say absence of a factor increase

OpenStudy (thomas5267):

\[ \begin{align*} &\quad(5\sqrt{5}+11)^{2n+1}\\ &=\sum_{k=0}^{2n+1}(5\sqrt{5})^k11^{2n+1-k}\\ &=\underbrace{\sum_{k=0}^n(5\sqrt{5})^{2k}11^{2n+1-2k}}_\text{integral part}+\underbrace{\sum_{k=0}^{n}(5\sqrt{5})^{2k+1}11^{2n-2k}}_\text{non-integral part} \text{ not sure did I get this right}\\ &=11\sum_{k=0}^n125^k121^{n-k}+\sum_{k=0}^{n}(5\sqrt{5})^{2k+1}11^{2n-2k}\\ &=\underbrace{11(246)^n}_\text{even}+\sum_{k=0}^{n}(5\sqrt{5})^{2k+1}11^{2n-2k} \end{align*} \]

OpenStudy (thomas5267):

Oh crap I left out the binomial coefficient lol.

ganeshie8 (ganeshie8):

they dont matter :)

ganeshie8 (ganeshie8):

then @dan815 's argument has a mistake because there are exactly \(n+1\) integer terms in the expanded sum

OpenStudy (thomas5267):

Why they don't?

ganeshie8 (ganeshie8):

because you figured it out without using them :) once a number is even, any multiple of it is even too... so..

OpenStudy (dan815):